# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v2

@@ IV.1.1 Strict NE as ESS
Let $a_{ij}$’s, $b_{ij}$’s, $c_{ij}$’s and $d_{ij}$’s denote the elements of the payoff matrices ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$, respectively. Without any loss of generality, consider the state under investigation be $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ which is a fixed point of the set of equations given by Eqs. (34a) and (34b). On doing linear stability analysis about it yields a diagonal Jacobian matrix, and hence the diagonal elements can be easily read as the eigenvalues—say, $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are $a_{1m}+b_{1n}-a_{mm}-b_{mn},a_{2m}+b_{2n}-a_{mm}-b_{mn},\cdots,a_{m-1,m}+b_{m-1,n}-a_{mm}-b_{mn},c_{1m}+d_{1n}-c_{nm}-d_{nn},c_{2m}+d_{2n}-c_{nm}-d_{nn},\cdots,c_{n-1,m}+d_{n-1,n}-c_{nm}-d_{nn}$. Here, comma has been put between subscripts wherever it is needed for clarity. It is clear that all the eigenvalues must be negative if the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ is a strict NE; thus, linear stability is automatically implied in such cases.
Now we would like to figure out whether the strict NE is ESS. For this purpose, we write the arbitrary vector in the neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ as $({\mathbf{x}},{\mathbf{y}})=((\delta_{x_{1}},\cdots,\delta_{x_{m-1}},1-\delta_{x_{1}}-\cdots-\delta_{x_{m-1}}),(\delta_{y_{1}},\cdots,\delta_{y_{n-1}},1-\delta_{y_{1}}-\cdots-\delta_{y_{n-1}}))$, where the $\epsilon$’s are infinitesimal and non-negative. A glance at Definition 1d directs us to calculate following terms, which we do up to first order in $\delta$’s:
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{m1}-a_{mm})\delta_{x_{1}}+(a_{m2}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\delta_{x_{m-1}},$ (35a)
$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\delta_{y_{1}}+(b_{m2}-b_{mn})\delta_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\delta_{y_{n-1}},$ (35b)
@@
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=-\delta_{x_{1}}\lambda_{1}-\delta_{x_{2}}\lambda_{2}-\cdots-\delta_{x_{m-1}}\lambda_{m-1}.$ (36)
@@
$\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}$ $=-\delta_{y_{1}}\lambda^{\prime}_{1}-\delta_{y_{2}}\lambda^{\prime}_{2}-\cdots-\delta_{y_{n-1}}\lambda^{\prime}_{n-1}.$ (37)
We can see that the first $m-1$ eigenvalues, as the coefficient of the $\delta_{x_{i}}$’s, come with negative signs in Eq. (36). Similarly, last $n-1$ eigenvalues, as the coefficient of $\delta_{y_{j}}$’s, come with negative signs in Eq. (37). If the fixed point has to be ESS then for both equations, the RHS’s have to be positive in the entire neighbourhood. This requirement equivalently demands that all the eigenvalues be negative because all the $\delta$’s are independent of each other and all are not allowed to be simultaneously zero. This demand is automatically satisfied if the fixed point is strict NE.
@@
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}-\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}.$ (38)
@@
either $\text{either }-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}>0~{}~{}\text{or}~{}~{}-\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}>0,$ (39)
@@ IV.1.2 Non-strict NE as ESS
The equivalence between ESS’s under and-, sum- and or-definitions goes away if the ESS is non-strict NE. Following example highlights this.
Example 4 (see Fig. 1d): Non-strict pure NE as ESS under or-definition need not be ESS under and- and sum-definitions.
@@
${\sf A}=\begin{pmatrix}a_{1}&a_{2}\\ a_{1}&a_{3}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}d_{1}&d_{2}\\ d_{1}&d_{3}\end{pmatrix},$ (40)
where $a_{2}>a_{3}$, $b_{2}>b_{3}$, $c_{2}<c_{3}$ and $d_{2}>d_{3}$. We can see that $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is a non-strict NE whose neighbourhood state is given as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. To check whether it is an ESS we calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\epsilon\delta(b_{2}-b_{3})+\epsilon^{2}(a_{2}-a_{3})$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=\epsilon\delta(c_{2}-c_{3})+\delta^{2}(d_{2}-d_{3})$. It is an ESS under or-definition since at least one of them is positive for all allowed values of $\epsilon$ and $\delta$. However, it is not an ESS under and-definition since $c_{2}-c_{3}<0$; furthermore, if $|c_{2}-c_{3}|>|b_{2}-b_{3}|$, then it may not satisfy an ESS condition under sum-definition. The fixed point remains asymptotically stable as may be checked numerically, thus, validating Theorem $1^{\prime}$.
Before going to the case of mixed ESS, we would like to bring it to the fore that for Definitions 1d and 2d, there are further arguments based on Lyapunov function for them to obey Theorem 1 in case of pure ESS corresponds to an isolated fixed point. One can show using the same Lyapunov function given in Eq. (20), that satisfying Definition 1d or 2d imply asymptotic stability for the pure ESS (corner fixed point of the simplotope). However, the general proof that Definition 3d imply asymptotic stability for pure ESS is not so straightforward: A proof using center manifold theory is available in the literature Cressman (1996).
@@ IV.2 Mixed ESS
The introduction of intra-specific interactions leads to circumventing certain restrictions present in bimatrix games. Specifically now and- and sum-definitions can lead to mixed ESS and mixed ESS under or-definition can be locally asymptotically stable. While the former is illustrated in Example 5 below, the latter is easily seen on choosing $m=n=2$, in which case the form of internal fixed point, it it exists, is
@@
It is evident that now there is a possibility of focus or node types of fixed points under suitable conditions. For clear visualization, consider the example of asymmetric game given in (44), where change in the stability of the fixed point—due to addition of intra-specific interaction—is demonstrated through phase portraits as plotted in Fig. 1(c) and Fig. 1(f). Furthermore, Example 6 (see below) makes an important point clear that 2ESS may help find new evolutionarily important asymptotically stable states which may escape and- and sum-definitions.
Example 5 (see Fig. 1e): Mixed ESS is possible under and- and sum-definitions in asymmetric games.
@@
${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (43)
One can easily check that the internal fixed point (mixed NE) in this case is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$. Therefore the neighbourhood vector has the form—$(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}{2}/{9}+\delta_{x},{7}/{9}-\delta_{x}\end{pmatrix},\begin{pmatrix}{2}/{9}+\delta_{y},{7}/{9}-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small and can be positive or negative but they can’t be zero simultaneously. We calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}={3\delta_{x}^{2}}/{4}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}={3\delta_{y}^{2}}/{4}$. Both terms are positive for all infinitesimal values of $\delta_{x}$ and $\delta_{y}$, therefore it satisfies and-definition. For sum-definition one just need to add both the terms and get, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}={3(\delta_{x}^{2}+\delta_{y}^{2})}/{4}$, which means sum-definition is also satisfied. Fig. 1(e) shows that the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$ is asymptotically stable, confirming Theorem 1 for and- and sum-definitions.
Example 6 (see Fig. 1f): Mixed 2ESS need not be ESS under and- and sum-definitions.
@@
${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}1&4\\ 2&2\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}6&4\\ 3&8\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (44)
Now, let see whether the internal fixed point is an ESS. One can calculate the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}66/115,49/115\end{pmatrix},\begin{pmatrix}238/345,107/345\end{pmatrix}\right)$, the neighbourhood vector therefore is, $(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}\frac{66}{115}+\delta_{x},\frac{49}{115}-\delta_{x}\end{pmatrix},\begin{pmatrix}\frac{238}{345}+\delta_{y},\frac{238}{345}-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small. We calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=3\delta_{x}\delta_{y}+\frac{3\delta_{x}^{2}}{4}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-7\delta_{x}\delta_{y}+\frac{3\delta_{y}^{2}}{4}$. It is not very hard observe that at least one of the terms in the RHS of the two expressions is positive for all possible infinitesimal values of $\delta_{x}$ and $\delta_{y}$, which means the internal fixed point is 2ESS—an ESS under or-definition. Since both the terms cannot remain simultaneously positive in the entire neighbourhood, it is definitely not an ESS under and-definition. To check for sum-definition we add the two terms and get, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-4\delta_{x}\delta_{y}+\frac{3(\delta_{x}^{2}+\delta_{y}^{2})}{4}$, which is not guaranteed to be always positive in the neighbourhood; hence it is not an ESS under sum-definition as well. Fig. 1(f) shows that the fixed point, a 2ESS, is asymptotically stable, confirming Theorem 1 for or-definition.
The stability of the mixed ESS can be understood through Lyapunov function. Let us consider the general case of mixed ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ when $m$ and $n$ are arbitrary by considering the following function given in Eq. (20). We calculate the time derivative using Eq. (34) and get,
@@
We can see that $\dot{V_{1}}<0$ $\forall({\mathbf{x}},{\mathbf{y}})$ in some neighbourhood of $({\mathbf{\hat{x}}},{\mathbf{\hat{y}}})$ when either inequalities (31) or inequalities (32) are satisfied. Hence, any mixed ESS under either and-definition or sum-definition has $V_{1}$ as a Lyapunov function and the corresponding fixed point is locally asymptotically stable. Note that, these conclusions does not carry over to bimatrix games (${\sf A}={\sf D}={\sf 0}$) for mixed ESS since we know only pure ESS is possible there. As far as the or-definition is concerned, a theorem Cressman (1996) comes to our help:
@@
qualifies as the Lyapunov function for a mixed 2ESS which, thus, is locally asymptotically stable fixed point. There is a subtlety to be noted: This existence of this Lyapunov function does not extend to bimatrix games (see Appendix A) formed by setting ${\sf A}={\sf 0}$ and ${\sf D}={\sf 0}$.

### v3

@@ IV.1.1 Strict NE as ESS
Let $a_{ij}$’s, $b_{ij}$’s, $c_{ij}$’s and $d_{ij}$’s denote the elements of the payoff matrices ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$, respectively. Without any loss of generality, consider the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ which is a fixed point of the set of equations given by Eqs. (31a) and (31b). On doing linear stability analysis about it yields a diagonal Jacobian matrix, and hence the diagonal elements can be easily read as the eigenvalues—say, $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are $a_{1m}+b_{1n}-a_{mm}-b_{mn},a_{2m}+b_{2n}-a_{mm}-b_{mn},\cdots,a_{m-1,m}+b_{m-1,n}-a_{mm}-b_{mn},c_{1m}+d_{1n}-c_{nm}-d_{nn},c_{2m}+d_{2n}-c_{nm}-d_{nn},\cdots,c_{n-1,m}+d_{n-1,n}-c_{nm}-d_{nn}$. Here, comma has been put between subscripts wherever it is needed for clarity. It is clear that all the eigenvalues must be negative if the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ is a strict NE; thus, linear stability is automatically implied in such cases.
Now we would like to figure out whether the strict NE is an ESS. For this purpose, we write the arbitrary vector in the neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ as $({\mathbf{x}},{\mathbf{y}})=((\delta_{x_{1}},\cdots,\delta_{x_{m-1}},1-\delta_{x_{1}}-\cdots-\delta_{x_{m-1}}),(\delta_{y_{1}},\cdots,\delta_{y_{n-1}},1-\delta_{y_{1}}-\cdots-\delta_{y_{n-1}}))$, where the $\epsilon$’s are infinitesimal and non-negative. A glance at Definition 1d directs us to calculate the following terms, which we do up to first order in $\delta$’s:
[…]
@@
[…]
@@
[…]
We can see that the first $m-1$ eigenvalues, as the coefficient of the $\delta_{x_{i}}$’s, come with negative signs in Eq. (33). Similarly, the last $n-1$ eigenvalues, as the coefficient of $\delta_{y_{j}}$’s, come with negative signs in Eq. (34). If the fixed point is an ESS under and-definition then for both equations, the RHS’s have to be positive in the entire neighbourhood. This requirement equivalently demands that all the eigenvalues be negative because all the $\delta$’s are independent of each other and all are not allowed to be simultaneously zero. This demand is automatically satisfied if the fixed point is strict NE.
@@
[…]
@@
$\text{either }-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}>0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ -\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}>0,$ (36)
@@ IV.1.2 Non-strict NE as ESS
The equivalence between ESSes under and-, sum- and or-definitions goes away if the ESS is non-strict NE. Following example highlights this.
Example 3 (see Fig. 1c): Non-strict pure NE as ESS under or-definition need not be ESS under and- and sum-definitions.
@@
${\sf A}=\begin{pmatrix}a_{1}&a_{2}\\ a_{1}&a_{3}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf D}=\begin{pmatrix}d_{1}&d_{2}\\ d_{1}&d_{3}\end{pmatrix},$ (37)
where $a_{2}>a_{3}$, $b_{2}>b_{3}$, $c_{2}<c_{3}$ and $d_{2}>d_{3}$. We can see that $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is a non-strict NE whose neighbourhood state is given as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. To check whether it is an ESS we calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\epsilon\delta(b_{2}-b_{3})+\epsilon^{2}(a_{2}-a_{3})$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=\epsilon\delta(c_{2}-c_{3})+\delta^{2}(d_{2}-d_{3})$. It is an ESS under or-definition since at least one of them is positive for all allowed values of $\epsilon$ and $\delta$. However, it is not an ESS under and-definition since $c_{2}-c_{3}<0$; furthermore, if $|c_{2}-c_{3}|>|b_{2}-b_{3}|$, then it may not satisfy the ESS condition under sum-definition. The fixed point remains asymptotically stable as may be checked numerically, thus, it satisfies Proposition 1.
Before going to the case of mixed ESS, we would like to bring it to the fore that, for Definitions 1d and 2d, there are further arguments based on Lyapunov function for them to obey Proposition 1 in case of pure ESS corresponding to an isolated fixed point. One can show using the same Lyapunov function given in Eq. (20), that satisfying Definition 1d or 2d imply asymptotic stability for the pure ESS (corner fixed point of the simplotope). However, the general proof that Definition 3d imply asymptotic stability for pure ESS is not so straightforward: A proof using center manifold theory is available in the literature Cressman (1996).
@@ IV.2 Mixed ESS
The introduction of intra-specific interactions leads to circumventing certain restrictions present in bimatrix games. Specifically now and- and sum-definitions can lead to mixed ESS (see Example 4) and mixed ESS under or-definition can be locally asymptotically stable (see Example 5). While the former is illustrated in Example 4 below, the latter is easily seen on choosing $m=n=2$, in which case the form of internal fixed point, if it exists, is
@@
It is evident that now there is a possibility of focus or node types of fixed points under suitable conditions. For clear visualization, consider the example of asymmetric game given in (41), where change in the stability of the fixed point—due to addition of intra-specific interaction—is demonstrated through phase portraits as plotted in Fig. 1(b) and Fig. 1(e). Furthermore, Example 5 (see below) makes an important point clear that 2ESS may help find new evolutionarily important asymptotically stable states which may escape and- and sum-definitions.
Example 4 (see Fig. 1d): Mixed ESS is possible under and- and sum-definitions in asymmetric games.
@@
${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf B}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf C}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (40)
One can easily check that the internal fixed point (mixed NE) in this case is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$. Therefore the neighbourhood vector has the form—$(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}{2}/{9}+\delta_{x},{7}/{9}-\delta_{x}\end{pmatrix},\begin{pmatrix}{2}/{9}+\delta_{y},{7}/{9}-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small and can be positive or negative but they can’t be zero simultaneously. We calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}={3\delta_{x}^{2}}/{4}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}={3\delta_{y}^{2}}/{4}$. Both terms are positive for all infinitesimal values of $\delta_{x}$ and $\delta_{y}$, therefore it satisfies and-definition. For sum-definition one just need to add both the terms and get, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}={3(\delta_{x}^{2}+\delta_{y}^{2})}/{4}$, which means sum-definition is also satisfied. Fig. 1(d) shows that the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$ is asymptotically stable, confirming Proposition 1 for and- and sum-definitions.
Example 5 (see Fig. 1e): Mixed 2ESS need not be ESS under and- and sum-definitions.
@@
${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf B}=\begin{pmatrix}1&4\\ 2&2\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf C}=\begin{pmatrix}6&4\\ 3&8\end{pmatrix}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (41)
Now, let see whether the internal fixed point is an ESS. One can calculate the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}66/115,49/115\end{pmatrix},\begin{pmatrix}238/345,107/345\end{pmatrix}\right)$, the neighbourhood vector therefore is, $(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}\frac{66}{115}+\delta_{x},\frac{49}{115}-\delta_{x}\end{pmatrix},\begin{pmatrix}\frac{238}{345}+\delta_{y},\frac{238}{345}-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small. We calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=3\delta_{x}\delta_{y}+\frac{3\delta_{x}^{2}}{4}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-7\delta_{x}\delta_{y}+\frac{3\delta_{y}^{2}}{4}$. It is not very hard observe that at least one of the terms in the RHS of the two expressions is positive for all possible infinitesimal values of $\delta_{x}$ and $\delta_{y}$, which means the internal fixed point is 2ESS—an ESS under or-definition. Since both the terms cannot remain simultaneously positive in the entire neighbourhood, it is definitely not an ESS under and-definition. To check for sum-definition we add the two terms and get, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-4\delta_{x}\delta_{y}+\frac{3(\delta_{x}^{2}+\delta_{y}^{2})}{4}$, which is not guaranteed to be always positive in the neighbourhood; hence it is not an ESS under sum-definition as well. Fig. 1(e) shows that the fixed point, a 2ESS, is asymptotically stable, confirming Proposition 1 for or-definition.
The stability of the mixed ESS can be understood through Lyapunov function. Let us consider the general case of mixed ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ when $m$ and $n$ are arbitrary, using the function given in Eq. (20). We calculate the time derivative using Eq. (31) and get,
@@
We can see that $\dot{V_{1}}<0$ $\forall({\mathbf{x}},{\mathbf{y}})$ in some neighbourhood of $({\mathbf{\hat{x}}},{\mathbf{\hat{y}}})$ when either inequalities (28) or inequalities (29) are satisfied. Hence, any mixed ESS under either and-definition or sum-definition has $V_{1}$ as a Lyapunov function and the corresponding fixed point is locally asymptotically stable. Note that, this conclusion for mixed ESS does not carry over to bimatrix games (${\sf A}={\sf D}={\sf 0}$) since we know only pure ESS is possible there. As far as the or-definition is concerned Theorem 1 (see the lemma in Appendix A of Ref. Cressman (1996) and the supplementary material of Ref. Cressman and Tao (2014)) below comes to our help:
@@
qualifies as the Lyapunov function for a mixed 2ESS which, thus, is locally asymptotically stable fixed point. There is a subtlety to be noted: This existence of this Lyapunov function does not extend to bimatrix games (see Appendix B) formed by setting ${\sf A}={\sf 0}$ and ${\sf D}={\sf 0}$.